Question 5 Answer the following questions What is the value of y when x is 15? y-10 + 2x If y is 10 what value is x? y = 10 + 2x What is the value of y when x is 10 and z is 2? y = 10+2x-2c what is the value of y when x is 0? y = 10 + 2x-2z What is the value of y when x is -5? y 10-6x +5z r
Answer #1

1. When x=15

y=10 + 2x

y=10 + 2(15)

y=10 +30 =40.

2. Y=10

10=10+2x

10-10=2x

2x=0

x=0

3. x=10 Z=2

y=10 + 2x - 2z

y= 10+2(10) -2(2)

y=10 +20-4

y=26

4. when x=0

y=10 +2(0) -2z

y=10+0-2z

y=10-2z

5 when x=-5

y=10-6(-5) +5z*-5

y=10 -6(-5) -25z

y=10+30 -25z

y=40-25z

-6 *-5=*30